PREVIOUS YEAR SOLVED PAPERS MCQ - GATE 2018

Correct AnswerOption A
Rate difficulty Rate difficulty:
Choose a difficulty.

Node that receives the packet to transmit will carrier-sense the medium for 5 units. Any packet which arrives within 5 unit will be sensed and keep the channel busy. Given that Signal speed is 10 meter/time.
Which means in 5 unit of time packet can travel 50 meters in max, that allow Q to successfully avoid the collision.
So, option (A) is correct.
Cancel reply

Your email address will not be published. Required fields are marked *


Cancel reply

Your email address will not be published. Required fields are marked *

Correct AnswerOption C
Rate difficulty Rate difficulty:
Choose a difficulty.
Item No.WeightValueValue/Weight
110606
27284
34205
422412
After Sorting
Item No.WeightValueValue/Weight
122412
210606
34205
47284
First, we have to Pick item_1 (Value weight ratio is highest).
After that Second highest is item_1. but cannot be picked because of its weight.
Now item_3 shall be picked. item_2 cannot be included because of its weight.
Therefore, overall profit by Vgreedy = 20+24 = 44
Hence, Vopt – Vgreedy = 60-44 = 16
So, answer is 16.
Cancel reply

Your email address will not be published. Required fields are marked *


Cancel reply

Your email address will not be published. Required fields are marked *

Correct AnswerOption A
Rate difficulty Rate difficulty:
Choose a difficulty.

Count(1024, 1024) will be called and value of x will be deducted by x/2. ‘x’ will be printed 10 times. On count=10, value of x will become 1.
For every recursively for every y=1023 and count() is called 10times for each y. So, 1023 x 10 = 10230
Answer is 10230.
Cancel reply

Your email address will not be published. Required fields are marked *


Cancel reply

Your email address will not be published. Required fields are marked *

Correct AnswerOption A
Rate difficulty Rate difficulty:
Choose a difficulty.
Share this question
Share:
The condition probability table given is
 HDMDLD
HG0.400.480.12
MG0.100.650.25
LG0.010.500.49

P(HG) = 0.2
P(MG) = 0.5
P(LG) = 0.3
Drawing the tree diagram for HD we get,

From diagram, P(HG ∩ HD) = 0.2x0.4
P(HD) = 0.2x0.4+0.5x0.1+0.3x0.01 = 0.133
Required probability,
P(HG | HD) = (0.2x0.4)/0.133 = 0.60
Cancel reply

Your email address will not be published. Required fields are marked *


Cancel reply

Your email address will not be published. Required fields are marked *

Correct AnswerOption C
Rate difficulty Rate difficulty:
Choose a difficulty.

Cancel reply

Your email address will not be published. Required fields are marked *


Cancel reply

Your email address will not be published. Required fields are marked *

Why Practice with Avatto?
High Quality MCQsExam-focused questions with accurate answers
Detailed ExplanationsStep-by-step explanations to enhance learning
Latest Exam PatternBased on updated syllabus and pattern
Performance AnalyticsTrack performance and improve weak areas
100% Free PracticePractice unlimited questions for free
Boost Your PreparationAttempt mock tests, previous year papers and topic-wise quizzes.Explore All Tests →